Given by the equations

                             (1)         Qd = 4 - P

                             (2)         Qs = -2+P           

a)    What is the price elasticity of the demand curve at the point of equilibrium?

b)    What is the price elasticity of the supply curve at the point of equilibrium?

1
Expert's answer
2021-07-29T12:26:04-0400

At equilibrium

"Qd=Qs\\4-P=-2+P\\4+2=P+P\\6=2P\\P=\frac{6}{2}\\P=3"


"Qd=4-P\\=4-3=1\\Qs=-2+P\\=-2+3=1"


At equilibrium:

P=3

Q=1


a)

Price elasticity of demand at equilibrium

"Ped=\frac{\Delta Q}{\Delta P}\times\frac{P}{Q}"


"\frac{\Delta Q}{\delta P}=-1"


"Ped=-1\times\frac{3}{1}"


"Ped=-3"


b)

Price elasticity of supply at equilibrium

"Es=\frac{\Delta Q}{\Delta P}\times \frac{P}{Q}\\"


"\frac{\Delta Q}{\delta P}=1"


"Es=1\times \frac{3}{1}"


"Es=3"